# Differentiable Fuzzy Cosmic-Web for Field Level Inference

### v1

@@ Abstract
Context. A comprehensive analysis of the cosmological large-scale structure derived from galaxy surveys involves field-level inference, which requires a forward modelling framework that simultaneously accounts for structure formation and tracer bias.
Aims. While structure formation models are well-understood, the development of an effective field-level bias model remains imprecise, particularly in the context of tracer perturbation theory within Bayesian reconstruction methods, which we address in this work.
Methods. To bridge this gap, we have developed a differentiable model that integrates augmented Lagrangian perturbation theory, nonlinear, nonlocal, and stochastic biasing. At the core of our approach is the Hierarchical Cosmic-Web Biasing Nonlocal (HICOBIAN) model, which provides a positive definite description of tracer bias while incorporating a long- and short-range nonlocal framework via cosmic-web regions and deviations from Poissonity in the likelihood. A key insight of our model is that transitions between cosmic-web regions are inherently smooth, which we implement using sigmoid-based gradient operations. This enables a fuzzy, and, hence, differentiable hierarchical cosmic-web description, making the model well-suited for machine learning frameworks.
Results. We demonstrate the efficiency of this model through GPU-accelerated computations implemented in JAX, the BRIDGE code, enabling scalable evaluation of complex biasing. Our approach accurately reproduces the primordial density field within associated error bars derived from Bayesian posterior sampling, as validated by two- and three-point statistics in Fourier space. Furthermore, we demonstrate that the methodology approaches the maximum encoded information consistent with Poisson noise. We also demonstrate that the bias parameters of a higher-order nonlocal bias model can be accurately reconstructed within the Bayesian framework.
Conclusions. We introduce a Bayesian field-level inference algorithm that leverages the same forward modelling framework used in galaxy, quasar, and Lyman alpha forest mock catalog generation– including nonlinear, nonlocal and stochastic bias with redshift space distortions–providing a unified and consistent approach to the analysis of large-scale cosmic structure.
@@ 1 Introduction
The primordial density fluctuations encode the compressed information of the entire cosmic history of our Universe. These early inhomogeneities, seeded during the initial moments after the Big Bang, evolved under gravity to form the vast cosmic-web structure we observe today. Reconstructing the initial conditions from present-day observations of luminous tracers–such as galaxies, quasars, Lyman-$\alpha$ forests, intensity maps, or from the 21cm line–offers a powerful probe of fundamental physics, including inflation, dark matter, dark energy, and gravity itself. The scientific community is making significant efforts to map the matter distribution in the Universe, as exemplified by current large-scale spectroscopic surveys such as DESI (Levi et al., 2013), Euclid (Amendola et al., 2016); and future ones such as PFS (Takada et al., 2014), MUST (Zhao et al., 2024) or ROMAN (Wang et al., 2022).
The first attempts to recover the primordial density fluctuations from the galaxy distribution were based on inverse reconstruction methods (e.g., Bertschinger & Dekel, 1989; Nusser & Dekel, 1992; Monaco & Efstathiou, 1999). These approaches proved highly effective for sharpening features such as the baryon acoustic oscillations (BAO) (Eisenstein et al., 2007), and were even proposed as tools to constrain primordial non-Gaussianities (Shirasaki et al., 2021).
@@
However, the accuracy of these inverse techniques is fundamentally limited by the non-invertibility of the gravitational evolution in the nonlinear regime in the absence of full phase-space information, particularly due to shell-crossing. Once multiple streams of matter overlap, information about the initial conditions is irreversibly lost, as the full phase-space dynamics–crucial for a unique reconstruction–are no longer accessible through the evolved tracer distribution alone.
@@
To overcome the limitations of inverse methods forward modelling was introduced as a principled alternative. Instead of inverting the non-linear structure formation process, forward approaches start from initial conditions and evolve them through a physical model of gravitational dynamics to generate predictions for present-day observables (Kitaura, 2013; Jasche & Wandelt, 2013; Wang et al., 2013; Jasche et al., 2015; Lavaux & Jasche, 2016). The first application of a forward modelling reconstruction to observational data was presented by Kitaura et al. (2012b). All these works demonstrated the viability of sampling high-dimensional posterior distributions consistent with both data and gravitational physics (including peculiar motions), laying the foundation for later advances in bias modelling and inference methods. Subsequent developments improved gravity solvers using particle-mesh codes (Wang et al., 2014; Jasche & Lavaux, 2019; Horowitz et al., 2019b) and emulators (Doeser et al., 2024). Further extensions incorporate light-cone evolution (Kitaura et al., 2021; Lavaux et al., 2019), primordial non-Gaussianities (Andrews et al., 2023), Effective Field Theory (EFT) approaches (Stadler et al., 2023), and hydrodynamics (Horowitz & Lukic, 2025) These techniques have also been applied to alternative tracers, such as the Lyman-$\alpha$ forest (e.g., Kitaura et al., 2012c; Horowitz et al., 2019a; Porqueres et al., 2019; Horowitz et al., 2021, 2022).
@@
This approach, however, faces a significant computational barrier. High-fidelity dark matter simulations capable of resolving the halos that host galaxies require substantial computational resources–often hundreds of thousands to millions of CPU hours–rendering brute-force posterior sampling impractical (see state-of-the-art $N$-body simulations, e.g., Garrison et al., 2018; Chuang et al., 2019; Ishiyama et al., 2021).
@@
To mitigate this challenge, Lagrangian perturbation theory (LPT) offers a computationally efficient alternative to full gravity solvers (Bernardeau et al., 2002), such as $N$-body simulations (Angulo & Hahn, 2022). While being approximate, LPT accurately captures gravitational evolution of matter on large, quasi-linear scales and serves as an effective forward model. Additional advancements have further extended and regularized these approximations, enabling reliable predictions down to Mpc–or even sub-Mpc– scales (Kitaura & Heß, 2013; Kitaura et al., 2024).
@@
This strategy is particularly effective when the forward model is constrained to its domain of validity and coupled to the observed tracers through an effective field-level bias prescription. The core challenge then lies in constructing a robust and physically motivated bias model. A critical first step is to ensure that large-scale bias is accurately captured, which necessitates the inclusion of nonlinear and nonlocal contributions up to third order (McDonald & Roy, 2009).
@@
A longstanding limitation of perturbation-theory-based bias models is that truncation at a fixed order often introduces unphysical behavior (McDonald & Roy, 2009; Schmittfull et al., 2019; Werner & Porciani, 2020). Specifically, these models can yield non–positive-definite densities and exhibit oscillatory or noisy behavior when extended into the highly nonlinear regime. However, recognizing that nonlocal bias terms at different orders correspond to distinct morphological features of the cosmic-web opens the door to a more stable and flexible framework (Kitaura et al., 2022). For each morphological feature, we can then assume a local nonlinear bias model, tailored to the specific properties of that structure. In this way, the classification of the cosmic-web into distinct patterns effectively acts as a diagonalizing operation, isolating the dominant local contributions to the tracer-matter relationship within each structural environment. This enables the systematic inclusion of higher-order terms while preserving physical consistency and numerical stability down to small scales, making the framework well-suited for precision field-level inference. This ansatz provides a general and flexible framework applicable to a wide range of tracer populations, and has demonstrated unprecedented accuracy in field-level bias modelling of point-like tracers such as dark matter halos (Balaguera-Antolínez et al., 2023; Coloma-Nadal et al., 2024) and the baryonic components resolved in cosmological hydrodynamical simulations–including ionized gas, neutral hydrogen, and the Lyman-$\alpha$ forest (Sinigaglia et al., 2021, 2022, 2024b, 2024a).
@@
In this work, we introduce BRIDGE 111github.com/pererossello/bridge-repo, a differentiable, GPU-accelerated framework for field-level inference written in JAX (Bradbury et al., 2018). In section 2, we detail the technical implementation of the structure formation and bias models. Section  3 presents a series of computational tests demonstrating that the code successfully recovers both the primordial density field and the bias parameters of a long-range nonlocal bias model at resolutions of 5 and 10 $h^{-1}\,\mathrm{Mpc}$. Finally, we apply BRIDGE using a combined long- and short-range bias model, showing that the framework can robustly recover the primordial density field even in the presence of highly complex tracer bias.
@@
Figure 1: Schematic overview of the BRIDGE pipeline. The framework combines a scalable, differentiable structure formation model with a flexible field-level bias prescription, all implemented in a GPU-accelerated JAX environment. This enables efficient Bayesian inference of the primordial density field from tracer observations, with support for complex, nonlocal bias models and multi-resolution analysis. The modular design allows for seamless integration of physical models while maintaining end-to-end differentiability and high computational performance.
@@ 2.1 Field-Level Inference
Let us consider a galaxy distribution, mapped onto a comoving volume by converting redshifts to distances, as the observed tracer field. We denote by $\boldsymbol{n}^{*}\in\mathbb{N}^{N^{3}}$ the observed tracer intensity on a cubic, comoving grid consisting of $N^{3}$ cells. A differentiable forward model $\mathcal{M}$ maps a white-noise realization $\boldsymbol{\nu}$, which defines the initial density field through a linear power spectrum, together with tracer bias parameters $\boldsymbol{b}$ and cosmological parameters $\boldsymbol{\Omega}$ to the expected tracer field $\bar{\boldsymbol{n}}$:
@@
The forward model $\mathcal{M}$ is constructed as a composition of three sequential maps: $\mathcal{M}=\mathcal{M}_{\text{b}}\circ\mathcal{M}_{\Psi}\circ\mathcal{M}_{\delta},$ where:
  1. 1.
    $\mathcal{M}_{\delta}$ is a linear transformation that maps a white-noise realization $\boldsymbol{\nu}$ to the initial overdensity field $\boldsymbol{\delta}_{0}$ at high redshift, using the linear matter power spectrum (see Appendix A for details on the input field parametrization).
  2. 2.
    $\mathcal{M}_{\Psi}$ is the gravity solver that evolves the initial field $\boldsymbol{\delta}_{0}$ into the nonlinear dark matter over-density field $\boldsymbol{\delta}$ at a specified redshift using an efficient gravity solver.
  3. 3.
    $\mathcal{M}_{\text{b}}$ maps the evolved matter field $\boldsymbol{\delta}$, together with the bias parameters $\boldsymbol{b}$, to the mean tracer field $\bar{\boldsymbol{n}}$, employing a cosmic-web-dependent, nonlocal bias model.
For each cell we assume the data arise from an independent discrete distribution with mean $\bar{n}_{i}$, which may include additional parameters $\boldsymbol{p}$ modelling survey completeness or super-Poissonian dispersion. The likelihood, which incorporates the forward model, reads
@@
where $\mathcal{P}(\boldsymbol{\nu})$ encodes prior information on the initial conditions, and $\mathcal{P}(\boldsymbol{b},\boldsymbol{p},\boldsymbol{\Omega})$ represents prior knowledge on the bias and cosmological parameters. The forward model is constructed such that the latent field $\boldsymbol{\nu}$ is whitened, therefore $\mathcal{P}(\boldsymbol{\nu})$ is an $N^{3}$-dimensional multivariate Gaussian with zero mean and identity covariance. The bias and additional likelihood model parameters, $\boldsymbol{b}$ and $\boldsymbol{p}$, which in our case will be strictly positive, are assigned independent log-normal priors. The probability density function of the log-normal distribution is given by:
$f(b;\,\mu,\sigma)=\frac{1}{b\sigma\sqrt{2\pi}}\exp\left[-\frac{(\ln b-\mu)^{2}}{2\sigma^{2}}\right]\,,\quad b>0\,,$ (4)
where $\mu$ and $\sigma$ are fixed hyperparameters that define the mean and standard deviation of $\ln b$. We adopt broad, uninformative values for these hyperparameters to allow the data to constrain the parameters effectively.
Throughout the remainder of this work, we fix the cosmological parameters $\boldsymbol{\Omega}$ to their fiducial values and omit their explicit dependence for notational simplicity.
@@ 2.2 Posterior Sampling
We exploit the differentiability of the forward model to efficiently sample the high-dimensional posterior using gradient-based Hamiltonian Monte Carlo (HMC) sampling (Duane et al., 1987). HMC was first brought to Bayesian large-scale-structure inference by Jasche & Kitaura (2010) and subsequently applied to observational data by Jasche et al. (2010).
We explore the posterior with HMC with the No-U-Turn Sampler (NUTS) in NumPyro (Hoffman & Gelman, 2011; Phan et al., 2019) and use the leapfrog integrator. During the adaptation (burn-in) phase the algorithm tunes the step size, trajectory length, and the mass matrix, which is kept diagonal due to memory constraints. During sampling, we fix the adapted mass matrix and step size and integrate the trajectories with $1024$ integration steps, resulting in the same number of gradient evaluations per sample.
To accelerate convergence each chain is initialized from a deterministic state obtained by Wiener-filtering the observed tracer density contrast (Zaroubi et al., 1995), $\boldsymbol{\delta}^{*}_{\text{tr}}\equiv\boldsymbol{n}^{*}N^{3}/N_{\text{tr}}-1$, back to an approximate initial density field $\boldsymbol{\delta}^{(0)}_{0}$, which is transformed into white noise $\boldsymbol{\nu}^{(0)}$.

### v3

# Differentiable fuzzy cosmic web for field-level inference

@@ Abstract
Context. A comprehensive analysis of the cosmological large-scale structure derived from galaxy surveys involves field-level inference, which requires a forward-modelling framework that simultaneously accounts for structure formation and tracer bias.
Aims. While structure formation models are well understood, the development of an effective field-level bias model remains challenging, particularly in the context of tracer perturbation theory within Bayesian reconstruction methods, which we address in this work.
Methods. To bridge this gap, we developed a differentiable model that integrates augmented Lagrangian perturbation theory and non-linear, non-local, and stochastic biasing. At the core of our approach is the Hierarchical Cosmic-Web Biasing Nonlocal (HICOBIAN) model, which provides a description of a field with a positive number of tracers while incorporating a long- and short-range non-local framework via cosmic-web regions and deviations from Poissonity in the likelihood. A key insight of our model is that transitions between cosmic-web regions are inherently smooth, which we implemented using sigmoid-based gradient operations. This enables a fuzzy and, hence, differentiable hierarchical cosmic-web description, making the model well-suited for machine-learning frameworks.
Results. We tested the practical implementation of this model through GPU-accelerated computations implemented in JAX, the BRIDGE code, enabling a scalable evaluation of complex biasing. Our approach accurately reproduces the primordial density field within associated error bars derived from Bayesian posterior sampling within a self-specified setting (meaning that inference is performed on data generated by the exact same forward model) as validated by two- and three-point statistics in Fourier space. Furthermore, we demonstrate that the methodology approaches the maximum encoded information consistent with Poisson noise. We also demonstrate that the bias parameters of a higher order non-local-bias model can be accurately reconstructed within the Bayesian framework for bias models with eight parameters.
Conclusions. We introduce a Bayesian field-level inference algorithm that leverages the same forward-modelling framework used in galaxy, quasar, and Lyman-alpha-forest mock-catalogue generation – including non-linear, non-local and stochastic bias with redshift space distortions – providing a unified and consistent approach to the analysis of large-scale cosmic structure.
@@ 1 Introduction
The primordial density fluctuations provide the initial conditions for structure formation and encode statistical signatures of the early Universe. These early inhomogeneities, seeded during the initial moments after the Big Bang, evolved under gravity to form the vast cosmic-web structure we observe today. Reconstructing the initial conditions from present-day observations of luminous tracers – such as galaxies, quasars, Lyman-$\alpha$ forests, intensity maps, or from the 21cm line – offers a powerful probe of fundamental physics, including inflation, dark matter, dark energy, and gravity itself. The scientific community is making significant efforts to map the matter distribution in the Universe, as exemplified by current large-scale spectroscopic surveys such as the Dark Energy Spectroscopic Instrument (DESI) (Levi et al., 2013; DESI Collaboration et al., 2025b), Euclid (Amendola et al., 2016; Euclid Collaboration et al., 2025); and future ones such as the Prime Focus Spectrograph (PFS) (Takada et al., 2014), MUltiplexed Survey Telescope (MUST) (Zhao et al., 2024), and the Nancy Grace Roman Space Telescope (Roman) (Wang et al., 2022).
@@
The first attempts to recover the primordial density fluctuations from the galaxy distribution were based on inverse reconstruction methods (e.g. Bertschinger & Dekel, 1989; Nusser & Dekel, 1992; Monaco & Efstathiou, 1999). These approaches proved to be highly effective for sharpening features such as baryon acoustic oscillations (BAOs; Eisenstein et al. 2007), and were even proposed as tools to constrain primordial non-Gaussianities (Shirasaki et al., 2021). However, the accuracy of these inverse techniques is fundamentally limited by the non-invertibility of the gravitational evolution in the non-linear regime in the absence of full phase-space information, particularly due to shell-crossing. Once multiple streams of matter overlap, information about the initial conditions is irreversibly lost, as the full phase-space dynamics – crucial for a unique reconstruction – are no longer accessible through the evolved tracer distribution alone.
@@
To overcome the limitations of inverse methods, forward modelling was introduced as a principled alternative. Instead of inverting the non-linear structure formation process, forward approaches start from initial conditions and evolve them through a physical model of gravitational dynamics to generate predictions for present-day observables (Kitaura, 2013; Jasche & Wandelt, 2013; Wang et al., 2013). The first application of a forward-modelling reconstruction (including peculiar motions) to observational data was presented by Kitaura et al. (2012b). All these works demonstrated the viability of sampling high-dimensional posterior distributions consistent with both data and gravitational physics, laying the foundation for later advances in bias modelling and inference methods. Subsequent developments considered improved mask handling (Jasche et al., 2015; Lavaux & Jasche, 2016; Ata et al., 2021), improved gravity solvers using particle-mesh codes (Wang et al., 2014; Jasche & Lavaux, 2019; Horowitz et al., 2019b), and emulators (Doeser et al., 2024). Further extensions incorporated light-cone evolution (Kitaura et al., 2021; Lavaux et al., 2019), primordial non-Gaussianities (Andrews et al., 2023), effective field theory (EFT) approaches (Stadler et al., 2023), and hydrodynamics (Horowitz & Lukic, 2025). These techniques have also been applied to alternative tracers, such as the Lyman-$\alpha$ forest (e.g. Kitaura et al., 2012c; Horowitz et al., 2019a; Porqueres et al., 2019; Horowitz et al., 2021, 2022). A benchmark of computational efficiency in this context is presented in Bayer et al. (2023).
@@
This approach, however, faces a significant computational barrier. High-fidelity dark-matter simulations capable of resolving the halos that host galaxies require substantial computational resources – often hundreds of thousands to millions of CPU hours – rendering brute-force posterior sampling impractical (see state-of-the-art $N$-body simulations, e.g. Garrison et al., 2018; Chuang et al., 2019; Ishiyama et al., 2021).
@@
To mitigate this challenge, Lagrangian perturbation theory (LPT) offers a computationally efficient alternative to full gravity solvers (Bernardeau et al., 2002), such as $N$-body simulations (Angulo & Hahn, 2022). While being approximate, LPT accurately captures the gravitational evolution of matter on large, quasi-linear scales and serves as an effective forward model. Additional advancements have further extended and regularised these approximations, enabling reliable predictions down to megaparsec – or even sub-megaparsec – scales (Kitaura & Heß, 2013; Kitaura et al., 2024).
@@
This strategy is particularly effective when the forward model is constrained to its domain of validity and coupled to the observed tracers through an effective field-level bias prescription. The core challenge then lies in constructing a robust and physically motivated bias model. A critical first step is to ensure that large-scale bias is accurately captured, which necessitates the inclusion of non-linear and non-local contributions up to the third order (McDonald & Roy, 2009).
@@
A longstanding limitation of perturbation-theory-based bias models is that truncation at a fixed order often introduces unphysical behaviour (McDonald & Roy, 2009; Schmittfull et al., 2019; Werner & Porciani, 2020). Specifically, these models can yield non-positive densities and exhibit oscillatory or noisy behaviour when extended into the highly non-linear regime. However, recognising that non-local-bias terms at different orders correspond to distinct morphological features of the cosmic web opens the door to a more stable and flexible framework (Kitaura et al., 2022). For each morphological feature, we can then assume a local non-linear-bias model tailored to the specific properties of that structure. In this way, the classification of the cosmic web into distinct patterns effectively acts as a diagonalising operation, isolating the dominant local contributions to the tracer-matter relationship within each structural environment. This enables the systematic inclusion of higher order terms while preserving physical consistency and numerical stability down to small scales, making the framework well-suited for precision field-level inference. This ansatz provides a general and flexible framework applicable to a wide range of tracer populations and has demonstrated unprecedented accuracy in field-level bias modelling of point-like tracers such as dark-matter halos (Balaguera-Antolínez et al., 2023; Coloma-Nadal et al., 2024) and the baryonic components resolved in cosmological hydrodynamical simulations – including ionised gas, neutral hydrogen, and the Lyman-$\alpha$ forest (Sinigaglia et al., 2021, 2022, 2024b, 2024a).
@@
In this paper, we introduce BRIDGE  [^1] Bayesian Reconstruction and Inference of Data-driven Generative Environments: github.com/pererossello/bridge-repo., a differentiable, GPU-accelerated framework for field-level inference written in JAX (Bradbury et al., 2018). In Sect. 2, we detail the technical implementation of the structure formation and bias models. Section 3 presents a series of computational tests demonstrating that the code successfully recovers both the primordial density field and the bias parameters of a long-range non-local-bias model at resolutions of 5 and 10 $h^{-1}\,\mathrm{Mpc}$. Finally, we describe how we applied BRIDGE using a combined long- and short-range bias model, showing that the framework can robustly recover the primordial density field even in the presence of highly complex tracer bias.
@@
Figure 1: Schematic overview of the BRIDGE pipeline. The framework combines a scalable, differentiable structure-formation model with a flexible field-level bias prescription, all implemented in a GPU-accelerated JAX environment. This enables efficient Bayesian inference of the primordial density field from tracer observations, with support for complex, non-local-bias models and multi-resolution analysis. The modular design allows for the seamless integration of physical models while maintaining end-to-end differentiability and high computational performance.
@@ 2.1 Field-level inference
We first consider a galaxy distribution, mapped onto a comoving volume by converting redshifts to distances, as the observed tracer field. With $\boldsymbol{n}^{*}\in\mathbb{N}^{N^{3}},$ we denote the observed tracer intensity on a cubic, comoving grid consisting of $N^{3}$ cells. A differentiable forward model, $\mathcal{M,}$ maps a white-noise realisation, $\boldsymbol{\nu}$, which defines the initial density field through a linear power spectrum, together with tracer bias parameters $\boldsymbol{b}$ and cosmological parameters $\boldsymbol{\Omega}$ to the expected tracer field $\bar{\boldsymbol{n}}$:
@@
The forward model, $\mathcal{M,}$ is constructed as a composition of three sequential maps, $\mathcal{M}=\mathcal{M}_{\text{b}}\circ\mathcal{M}_{\Psi}\circ\mathcal{M}_{\delta},$ where $\mathcal{M}_{\delta}$ is a linear transformation that maps a white-noise realisation, $\boldsymbol{\nu,}$ to the initial overdensity field, $\boldsymbol{\delta}_{0},$ at high redshift using the linear matter power spectrum (see Appendix A for details on the input field parametrisation); $\mathcal{M}_{\Psi}$ is the gravity solver that evolves the initial field, $\boldsymbol{\delta}_{0},$ into the non-linear dark-matter over-density field, $\boldsymbol{\delta,}$ at a specified redshift using an efficient gravity solver; and $\mathcal{M}_{\text{b}}$ maps the evolved matter field, $\boldsymbol{\delta}$, together with the bias parameters, $\boldsymbol{b}$, to the mean tracer field, $\bar{\boldsymbol{n}}$, employing a cosmic-web-dependent, non-local-bias model.
For each cell, we assume the data arise from an independent discrete distribution with mean $\bar{n}_{i}$, which may include additional parameters, $\boldsymbol{p,}$ modelling survey completeness or super-Poissonian dispersion. The likelihood, which incorporates the forward model, reads
@@
where $\mathcal{P}(\boldsymbol{\nu})$ encodes prior information on the initial conditions, and $\mathcal{P}(\boldsymbol{b},\boldsymbol{p},\boldsymbol{\Omega})$ represents prior knowledge on the bias and cosmological parameters. The forward model was constructed such that the latent field, $\boldsymbol{\nu,}$ is whitened, and therefore $\mathcal{P}(\boldsymbol{\nu})$ is an $N^{3}$-dimensional multi-variate Gaussian with zero mean and identity covariance. The bias and additional likelihood model parameters, $\boldsymbol{b}$ and $\boldsymbol{p}$, which in our case were strictly positive, were assigned independent log-normal priors. The probability density function of the log-normal distribution for the bias parameters implemented in the BRIDGE code is given by
$\boxed{f(b;\,\mu,\sigma)=\frac{1}{b\sigma\sqrt{2\pi}}\exp\left[-\frac{(\ln b-\mu)^{2}}{2\sigma^{2}}\right]\,,\quad b>0}\,$, (4)
where $\mu$ and $\sigma$ are fixed hyperparameters that define the mean and standard deviation of $\ln b$. Throughout the remainder of this work, we fixed the cosmological parameters, $\boldsymbol{\Omega,}$ to their fiducial values and omitted their explicit dependence for notational simplicity.
@@ 2.2 Posterior sampling
We exploited the differentiability of the forward model to efficiently sample the high-dimensional posterior using gradient-based Hamiltonian Monte Carlo (HMC) sampling (Duane et al., 1987). HMC was first applied to Bayesian large-scale-structure inference by Jasche & Kitaura (2010) and subsequently to observational data by Jasche et al. (2010).
We implemented within the BRIDGE code various posterior sampling methods, such as second- and fourth-order HMC sampling and the Blanes et al. (2014) HMC variant. The numerical tests shown in this work were finally performed with the publicly available No-U-Turn Sampler (NUTS) in NumPyro (Hoffman & Gelman, 2011; Phan et al., 2019). During the adaptation (burn-in) phase, the algorithm tunes the step size, trajectory length, and the mass matrix, which is kept diagonal due to memory constraints. During sampling, we fixed the adapted mass matrix and step size and integrated the trajectories with a NUTS tree depth equal to 10.
To accelerate convergence, each chain was initialised from a deterministic state obtained by Wiener-filtering the observed tracer density contrast (Zaroubi et al., 1995), $\boldsymbol{\delta}^{*}_{\text{tr}}\equiv\boldsymbol{n}^{*}N^{3}/N_{\text{tr}}-1$, back to an approximate initial density field, $\boldsymbol{\delta}^{(0)}_{0}$, which was transformed into white noise, $\boldsymbol{\nu}^{(0)}$; $N_{\rm tr}$ is the total number of tracers. In this first step, we assumed the linear power spectrum to be known and ignored peculiar motions.